wiki:Research/Instrumentation/Meetings/2019/09-18

SmartPhantom meeting: Thursday 29th August 2019, 11:00 BST

Present: KL, LC, JB, HTL, AK

Notes:

  1. Notes and actions from last meeting:
    • KL: Liaise to develop concept for polishing.
      • Stands: hold until polishing methodology is resolved.
    • AK: Prepare paperwork necessary to gain permission to work with source in 527.
      • Done.
    • KL: Check whether there was enough money to buy another 150mm stage.
      • Done. There was. Issue to be revisited today in the light of the requirement to purchase more fibre.
    • KL: Let D.Loader know that the other equipment may have to be sent without the SmartPhantom stuff.
      • Done.
  1. Baseline schedule and progress against schedule: KL/HTL
    The schedule was discussed and updated (see updated schedule). We discussed the schedule and agreed that we need to make a revision to reflect the present status of the development (action KL).

Points noted:

  • LabView readout is ready for the test of the SciWire. Important now to define how teh detector will be read out at MedAustron. This could be a video (60 frms/sec). Agreed to add action KL, AK, HTL to consult to specify the data recording at MedAustron while the tests on SciWire are being carried out.
  1. Experience with trial windings and discussion: LC, LTL
    • Detailed log of prototyping to date.
      • HTL presented the key points of the comprehensive note. The excellent discussion was hard to minute. However, we agreed that the next trial winding will be with the modified jig, circular aperture in the carbon-fibre frame, and perhaps test use of the various ideas put forward, including:
        • Clamping the fibure during winding;
        • Exploitation of adhesive tape to restrain fibre wile winding; and
        • Using adhesive spray to generate an adhesive surface to restrain the fibres during winding.
    • We agreed to "publish" HTL's note as a CCAP note. AK will invent a note numbering scheme which HTL will add to the header of the note so that he can post it as a CCAP note.
  1. Mechanical design/manufacture update: LC
    • Design update.
      • LC presented the proposed modifications to winding jig and carbon-fibre frame. In the discussion, JB proposed that the fibre be annealed post winding to relieve the tension. The operatng range of the fibre from the data sheet is -20C -- +50C. So, annealing will need to be done at a temperatrure >50C.
      • We agreed to investigate whether this has been done by any other groups.
      • Agreed to make next test with a circular apperture.
      • KL will email D.Nardini and I.Clark to explain the teh next steps in the programme and to request machining of carbon fibre (two frames, two connector strips).
  1. Control system and SciWire readout update: AK
    • Progress to date noted under schedule discussion.
    • Connection of cameras to LapTop requires USB hub. Need to purchase additional (action AK) cameras and hubs. LapTop has three USB ports, so need 3 USB hubs. We agreed to buy hard-drive to store data but to defer the camera purchase.
  1. Optical readout: JB
    • Optical system specified. Outstanding issue is the detailed design of the piece that connects the fibre to the focusing system. We agreed to call this the boot. We agreed to take the optical components to LC so that he can finalise the design of the boot (actions: transport of componenents KL/HTL, finalised design LC).
    • Need to specify boot -- LC needs to know what it connects to. Take focusing device to LC.
  1. Simulation/analysis update: HTL
    • HTL had re-evaluated the impact of using a singlet layer on the performance of the SmartPhantom. The results, consistent with the last pass of single-layer evaluation, indicated that the change had a modest impact on performance.
  1. Spend status, discussion of outstanding requirement: KL/All
    Spend to date.
    • We discussed the outstanding purchases that were required/being considered:
      • Scintilalting and clear fibre: estimate fibre cost for 5 stations: 150x150mm2! square, 2 doublets, five stations: £18k.
        • Actions: KL send fibre invoices to HTL; HTL provide spreadsheet with revised, precise estimate of fibre cost.
      • Carbon fibre for frames (may need to order more)
      • Material/components for mounting of stations;
      • Cameras and associated optics
      • USB hubs and hard disk
      • Fishing line for tests.
    • We noted that the following items were not included in the costing:
      • Camera £469.20
      • Optics £356.10
      • Locking rings (£30.00) were included in stepper motor orer.
    • KL will prepare a revised spreadsheet for next meeting.
  1. Next steps: All
    • We agreed:
      • New winding test with circular frame; implies need for new frames and connector pieces;
      • SciWire tests; this will be also a prototyping of the optical r/p for SmartPhantom.

  1. DONM
    • Meet Tuesday 10:00
  1. AoB
    • None.

Summary of actions:

  • KL: Liaise to develop concept for polishing.
  • KL: Prepare revision to schedule for the next meeting.
  • KL,AK,HTL: Consult to specify the data recording at MedAustron.
  • AK: Invent a note numbering scheme so that HTL's note can be "published" on the wiki as a CCAP note.
  • HTL: Add CCAP note number to the header of the note and post it as a CCAP note.
  • KL: Email D.Nardini and I.Clark to explain the teh next steps in the programme and to request machining of carbon fibre (two frames, two connector strips).
  • AK: Purchase hard-drive to store SmartPhantom data in operation.
  • KL or HTL: Transport optical components from Imperial to RAL (LC).
  • LC: Complete detailed design of the boot.
  • KL: Send fibre invoices to HTL.
  • HTL: Provide spreadsheet with revised, precise estimate of fibre cost.
  • KL: Prepare a revised costing spreadsheet for next meeting.

Meeting details

Venue: my apologies; I did not book a room at Imperial!

Perhaps JB, AK, and HTL et al can work out a convenient common location.

Phone details:

  UK: 08082 380274
  Austria: 0800 006612 

  PIN: 5904750902

Full list of telephone contacts may be found here.


Agenda:

  1. Notes and actions from last meeting:
    • KL: Liaise to develop concept for polishing.
      • Stands: hold until polishing methodology is resolved.
    • AK: Prepare paperwork necessary to gain permission to work with source in 527.
      • Done.
    • KL: Check whether there was enough money to buy another 150mm stage.
      • Done. There was. Issue to be revisited today in the light of the requirement to purchase more fibre.
    • KL: Let D.Loader know that the other equipment may have to be sent without the SmartPhantom stuff.
      • Done.
  1. Baseline schedule and progress against schedule: KL/HTL
    The schedule was discussed and updated (see updated schedule).
  1. Experience with trial windings and discussion: LC, LTL
  1. Mechanical design/manufacture update: LC
  1. Control system and SciWire readout update: AK
  1. Optical readout: JB
  1. Simulation/analysis update: HTL
  1. Spend status, discussion of outstanding requirement: KL/All
    Spend to date.
  1. Next steps: All

  1. DONM
  1. AoB

Last modified 2 years ago Last modified on Sep 19, 2019 5:52:53 AM

Attachments (4)

Download all attachments as: .zip